OpenStudy (anonymous):

Bill is selling stickers to raise money for a charity. He sold 36 stickers on the first day and 23 on the second day. He had 26 stickers left at the end of the second day. How many stickers did Bill have to start with?

OpenStudy (kropot72):

36 + 23 + 26 = 85

OpenStudy (albert0898):

First Day = 36 Stickers Second Day = 23 Stickers Remainder after Day Two = 26 Total Amount Before Any Sales = 36 + 23 + 26 = 85 Your final answer is 85

OpenStudy (anonymous):

Thank You !